A moving train travels through station P with a speed of 20 meter per seconds anf accelerate uniformly for 2 seconds reaching a speed of 48 meter per second, it continue with this speed for 5 seconds and then retardate uniformly for another 3 seconds and comes to rest at station Q. Find the distance from P to Q
1
Expert's answer
2020-09-09T10:20:18-0400

Solution:

acceleration :


a = "\tfrac{(V_{f} - V_{i})}{2}" = "\\tfrac{48-20}{2}\t=14 (\\tfrac{m}{s^{2}}\t)"


acceleration time :

ta = 2.0 sec


retardation :

r = "\\tfrac{0-48}{3}=-16 (\\tfrac{m}{s^{2}}\t)" = -16 m/sec^2


retardation time:

tr = 3.0 sec


distance computation :

"S_{1}=V_{i}\\cdot t_{a}\\cdot \\tfrac{a\\cdot t_{a}^{2}}{2}\t=68m"

"S_{2}=V_{f}\cdot 5=240m"

"S_{3}=V_{f}\\cdot \\tfrac{t_{r}}{2}\t=72m"

"S_{total}=S_{1}+S_{2}+S_{3}=380m"


The distance from P to Q is 380 m.